# Gas contamination and mitigation in a 100 m3 / 10 bar argon TPC with optical readout: a viability study

### v1

@@ 2.2 Outgassing model
Outgassing is dominated by the technical plastics employed in TPC construction, that is orders of magnitude above that of metals, e.g., from the TPC vessel or field cage. We consider Teflon (commonly used as a light reflector and as a structural material for field cages [33, 34]), PMMA (a sturdy material with good optical properties) and PEEK (a low-outgassing structural material). Plastics follow, in general, an outgassing law depending on time ($t$) as $Q(t)\sim t^{-a}$ with $a\sim 1/2$, corresponding to a diffusion-driven bulk desorption law. This is in principle applicable as long as i) the desorbed gas remains far from the saturation concentration of the exterior (about $f=3000$ ppm in case of H2O at ${}^{\circ}$, P=10 bar) and ii) the material is thick enough to keep furnishing the contaminant at any given time. As time passes, however, the desorbed gas reaches out from deeper regions until the material boundaries are approached, moment at which outgassing drops exponentially. The process leads to the following piece-wise outgassing law [35]:
$Q(t)=\begin{cases}k\sqrt{\frac{\pi\tau}{t}}\quad~{}~{}~{}~{}\text{for }t\leq 0.5\tau,\\ 4\,k\,e^{-t/\tau}\quad\text{for }t>0.5\tau.\\ \end{cases}$ (7)
@@
Here, $k$ indicates the strength of the outgassing rate and $\tau$ is the time constant that characterizes the moment at which the depth-of-origin of the desorbed species approaches the material thickness $t_{m}$. It is computed as $\tau=\frac{t_{m}^{2}}{\pi^{2}D_{\text{bulk}}}$, with $D_{\text{bulk}}$ being the diffusion coefficient in the bulk. A typical value for technical plastics is $D_{\text{bulk}}=(3.0\pm 1.2)\times 10^{-9}$ cm2 s-1, following [36].
@@
Outgassing rates are usually obtained in vacuum, and they can be parameterized through the value measured 10 hours after pumping ($Q_{10h}$) and the exponent $a$ [37] (the latter assumed to be 0.5 hereafter). For instance, in the case of Teflon at room temperature, $Q_{10h}$ varies typically from $3.5\times 10^{-8}$ to $1\times 10^{-7}\textnormal{mbar}\cdot{\textnormal{l}\cdot\textnormal{cm}^{-2}\cdot{\textnormal{s}^{-1}}}$ (for a recent review, see [38]). Mass spectroscopy indicates that the outgassing content in O2 is generally less than 10% of the total while N2 and H2O concentrations vary much more, going from nearly 70% H2O and 30% N2/CO to nearly 100% N2 and 0% H2O [19, 37, 39]. For both H2O and N2 we will consider the most unfavourable case (i.e., that they constitute 100% of the outgassing), together with the middle (‘typical’) and lowest (‘optimistic’) $Q_{10h}$ values obtained for the total outgassing rates in [38]. For O2 we take a range comprised between the value directly measured in [19] (‘optimistic’) and 10% of the middle value of $Q_{10h}$ in [38] (‘typical’). As outgassing rates depend on the material quality and surface finish, the rationale for our choice has been to select a range that includes typical outgassing values as well as the lowest ones found in literature. Measurements in [19] are important, as they do not involve bake out, thereby matching the expected experimental conditions under which large-volume TPCs are commissioned.
@@
Outgassing rates for PMMA and PEEK, on the other hand, are found in the range $Q_{10h}=10^{-7}$$10^{-6}$ $\frac{\textnormal{mbar}\cdot{\textnormal{l}}}{\textnormal{cm}^{2}\cdot{\textnormal{s}}}$. In the former case, they tend to be dominated by H2O, without any trace of O2 or N2 being found (e.g., [19]). In the latter, the situation is similar however the amount of O2 reported in literature is %-level (e.g. [40]). Assuming PMMA to be typically used as a window for light detection [41] and PEEK to be sometimes present in readout frames [42], one can anticipate that the use of these materials would be restricted to the end-caps, constituting a 20% of the total outgassing surface (at maximum) in our default geometry. Thus, the contribution of PMMA and PEEK to outgassing can be made generally subdominant in the presence of a Teflon reflector, if the former materials are chosen in the lower-end of the range of measured values.
@@
In the following, we will adopt the outgassing law in eq. 7 for studying the impact of H2O and N2 in the TPC performance. For O2 (a priori the most critical case), we will perform the calculations assuming a more extreme scenario in which the outgassing remains constant at $Q=Q_{\text{10h}}$. In all cases, Teflon will be assumed to dominate the system outgassing. A compilation of the outgassing rates considered in this work can be found in table 1 (column 2).
@@
$D_{\text{x, He/Ne/Ar/Xe}}=\frac{0.00143\cdot T^{1.75}}{P\sqrt{\frac{2}{\frac{1}{M_{\text{x}}}+\frac{1}{M_{\text{He/Ne/Ar/Xe}}}}}\cdot[\sqrt[3]{V_{\text{x}}}+\sqrt[3]{V_{\text{He/Ne/Ar/Xe}}}]^{2}},$ (8)
@@
$\frac{P}{D_{\text{x},B_{1}+B_{2}}}=\frac{P_{B_{1}}}{D_{\text{x},B_{1}}}+\frac{P_{B_{2}}}{D_{\text{x},B_{2}}}.$ (9)
@@
It should be noted that it is in principle possible that additional mechanisms at the material-gas interface could introduce a pressure dependence and an outgassing law different from the one assumed in vacuum. This would need to be elucidated through direct measurements. As for other assumptions employed earlier in text, the aim of our study is to provide upper limits to the impact of impurities in an actual system, and verify that they can be brought below potentially harmful levels through reasonable design choices. Along this direction, there is no obvious mechanism by which outgassing could be enhanced as pressure increases, compared to the model proposed here.
@@ 2.3 Gas distribution scheme
Our baseline proposal for the gas distribution is based on axial rods (perforated plastic tubes) attached to the field-cage, the inlets being at the bottom-half and the outlets at the top-half, as shown in Fig. 4 (top-left). The rationale for this choice is to allow a simple implementation since the field-cage is generally the less instrumented part of a TPC. Thus, this concept does not interfere with the end-cap design and, with sensible choices, it can assure uniform gas distribution along the drift/axial direction. Sixteen equispaced axial rods run along the rectangular section of the TPC (eight for injection and eight for extraction) as shown in Fig. 4 (top-right), providing an inter-rod distance of nearly 1 m between the centers. As shown later, this is sufficient to sweep the outgassing from the field-cage walls efficiently. The distributor rods are periodically perforated with a pattern that is discussed in the next section.
@@
It is assumed that the gas is continuously recirculated through compressors that, in combination with purifiers (getters), can keep the injected gas clean at all times. Commercial ‘cold’ getters are capable of maintaining O2 and H2O concentrations at ppb-level per pass, being at the same time oblivious to common molecular additives such as CH4 and CF4. In simulation, as the purification levels are at least two orders of magnitude below the concentrations considered to be harmful, we will assume that the injected gas is at all times 100% pure. On the other hand, getters capable of trapping N2 need of heating, that may cause incompatibility with high gas flows as well as with the use of molecular additives. When studying nitrogen contamination in text, we will not consider the possibility of purification.
Given the above conditions, and in the absence of gas leaks, the ultimate purity levels that can be achieved are limited by the power of the recirculation compressors. For reference, the ALICE TPC circulates gas at 20 m3/h (1 bar) while the NEXT-100 experiment is designed to circulate at around 5 m3/h (10 bar), [45]. Based on this, we will consider two different conditions for a 10 bar system, referring to them as ‘typical’ (30% above NEXT flow: 6.5 m3/h) and ‘optimistic’ (20 m3/h), as summarized in table 2.
@@ 3.1 Optimization of the distributor rods
Results from a three-dimensional CFD simulation performed, illustratively, for a 3-hole distributor, are shown in Fig. 5. It provides a cross-section at the holes’ middle plane (the modulus of the gas velocity is represented on the left column, the pressure profile on the right one). Two reference cases are considered: a capped distributor (top) and an open one (bottom). As expected, an open distributor is more inefficient, given that only a fraction of the gas exits the distributor through the holes and into the chamber, as small as few % for high gas flow. It also leads to a higher pressure drop along the distributor due to friction, a far greater directionality of the outgoing gas and a non-uniform discharge pattern. Low efficiency in combination with a directional and non-uniform discharge pattern will impact the distribution of impurities throughout the chamber and the time to achieve an uniform mixing with the main gas. Therefore, we opted for a capped distributor, assuring uniform and normal injection of all the circulated gas.
A large number of holes per gas distributor, of the order of several 100’s, has been proposed in previous large-volume TPCs (e.g. [11]). Optimizing such a geometry, each hole being potentially down to $\mathcal{O}($mm) in size, and each distributor rod extending over a length of 555~{}5m, is impractical when using numerical simulations. As an alternative, we opted for a parametric approach based on the balance equations for each hole section, that allows to obtain the solution iteratively. We followed the classical formulation of [48], whereby: i) a discharge coefficient $C_{d}$ (between 0 and 1) regulates the loss of flow into the distributor holes,  — mass balance; ii) a regain coefficient $C_{r}$ (between 0 and 2) regulates the pressure increase along the distributor after the ensuing velocity drop (Fig. 5-right),  — momentum balance; iii) an angular coefficient $\gamma$ regulates the angle of discharge relative to the normal:
$D^{2}(v_{\text{up}}-v_{\text{down}})$ $=$ $d^{2}C_{d}\sqrt{\frac{\Delta{P_{\text{hole}}}}{\rho/2}},$ (10)
$P_{\text{down}}-P_{\text{up}}$ $=$ $C_{r}\frac{1}{2}\rho(v_{\text{up}}^{2}-v_{\text{down}}^{2}),$ (11)
$\theta$ $=$ $\gamma\cdot\arctan\left(\sqrt{\frac{\frac{1}{2}\rho v_{\text{up}}^{2}}{\Delta{P}_{\text{hole}}}}\right).$ (12)
@@
$\Delta{P_{\text{hole}}}=\frac{P_{\text{up}}+P_{\text{down}}}{2}-P_{{}_{\text{TPC}}}.$ (13)
[…]
Figure 6: Main magnitudes characterizing the response of the distributor rod for each triple-hole along the distributor, shown with red lines (the first triple-hole upstream is number 1). We consider an argon gas flow of 6.5 m3/h at 10 bar, 8 distributors of 3 cm in diameter and 5 m in length. The hole size is 0.75 mm and holes are separated by 5 cm, (geometry $D$ in table 2). Top-left: gas velocity in the distributor at the hole position. Top-right: discharge velocity. Bottom-left: angle relative to normal. Bottom-right: pressure difference relative to the TPC. (The analytical expressions given in text are shown as dashed lines).
In the conditions studied here, as $d\ll{D}$ and the discharge velocity is both small and very similar for all holes, it is possible to approximate the above magnitudes by (see appendix):
$\Delta{P}_{\text{hole}}$ $\simeq$ $\frac{1}{2}\frac{1}{C_{d}^{2}}\rho v_{\text{hole}}^{2},$ (14)
$v_{\text{hole}}$ $\simeq$ $\frac{\Phi}{\textnormal{t.h.a.}},$ (15)
$\theta_{\text{max}}$ $\simeq$ $\gamma C_{d}\frac{d^{2}}{D^{2}}\frac{L_{\text{dist}}}{\Delta{L}},$ (16)
@@
These analytical estimates are included in Fig. 6 as blue dashed lines, showing an excellent agreement with the numerical solution. It should be noted that, with independence from the distributor geometry, $C_{d}$ approaches a constant value around $C_{d}=0.61$-$0.64$ when $\Delta{P}_{\text{hole}}\gg\frac{1}{2}\rho v_{\text{dist}}^{2}$ [49]. A detailed comparison with experimental data is presented in [48], providing an estimate of $C_{d}=0.63$ (that is adopted hereafter). Such pressure-dominated conditions are typical of capped distributors, and were verified for all our study cases. Similarly, the $\gamma$ pre-factor is largely independent from the gas flow characteristics and, according to [48], weakly dependent on the geometry, and so the value is set to $\gamma=0.71$. The least constrained coefficient, $C_{r}$, as well as the friction losses, are immaterial in the conditions discussed, as it is made clear through the analytic expresions in 14-16. The aforementioned values for the parameters $\gamma$ and $C_{d}$ are consistent with the CFD analysis of the control geometry displayed in Fig. 5.
@@
In order to come to some plausible distributor geometries, the following design constraints were introduced: i) a diameter hole not much smaller than 1 mm, ii) a diameter of the distributor not much larger than 5 cm, iii) an inter-hole distance not larger than 5 cm, iv) a pressure drop below 10 mbar, v) a discharge velocity well below the speed of sound, and vi) a maximum discharge angle below or around $5^{\circ}$. Based on these requirements, the main challenge is to obtain a small discharge angle without neither letting the inter-hole distance and tube diameter grow too big, nor the hole diameter become too small. Four exemplary geometries satisfying the above conditions ($A$-$D$) are compiled in table 2, and we have focused in the remainder of the text on $A$ (single hole) and $B$ (triple hole). It might be argued that small distributor diameters like those given in $C$ and $D$ should be preferred. This can be in fact achieved for $A$ and $B$ as well, through minor adjustments of the geometry that have not been considered in the CFD simulations presented in this work. For instance, a reduction from a diameter of 7 cm to 3.5 cm in geometry $B$ can be achieved by increasing the number of injection rods to 32 and increase the inter-hole distance to 60 mm, yet keeping the rest of the fluid properties as in table 2. Naturally, it is also possible to reduce the diameter to 3 cm or less and maintaining a small discharge angle through the increase in the thickness of the distributor walls, a study that falls outside the scope of this work. A slight preference towards geometries $A$ and $B$ stems from the fact that hole velocities above 1-2 m/s (as in $C$, $D$) showed a tendency to form large-scale vortexes, that delay substantially the time for the impurities to achieve uniform mixing. Aiming at a comparative study, we opted for a design that guaranteed a range of velocities in the default distributor geometry ($B$) outside the region where the effect appears, for the gas flows studied. A CAD image of this geometry is shown in Fig. 7-a, whereas Fig. 7-b provides an equivalent geometry in terms of fluid parameters at the inlets, however with smaller-diameter distributors.

### v2

@@ 2.2 Outgassing model
Outgassing is dominated by the technical plastics employed in TPC construction, that is orders of magnitude above that of metals, e.g., from the TPC vessel or field cage. We consider Teflon/PTFE (commonly used as a light reflector and as a structural material for field cages [35, 36]), PMMA (a sturdy material with good optical properties) and PEEK (a low-outgassing structural material). Plastics follow, in general, an outgassing law depending on time ($t$) as $Q(t)\sim t^{-a}$ with $a\sim 1/2$, corresponding to a diffusion-driven bulk desorption law. This is in principle applicable as long as i) the desorbed gas remains far from the saturation concentration of the exterior (about $f=3000$ ppm in case of H2O at ${}^{\circ}$, P=10 bar) and ii) the material is thick enough to keep furnishing the contaminant at any given time. As time passes, however, the desorbed gas reaches out from deeper regions until the material boundaries are approached, moment at which outgassing drops exponentially. The process leads to the following piece-wise outgassing law [37]:
$Q(t)=\begin{cases}k\sqrt{\frac{\pi\tau}{t}}\quad~{}~{}~{}~{}\text{for }t\leq 0.5\tau,\\ 4\,k\,e^{-t/\tau}\quad\text{for }t>0.5\tau.\\ \end{cases}$ (2.7)
@@
Here, $k$ indicates the strength of the outgassing rate, and $\tau$ is the time constant that characterizes the moment at which the depth-of-origin of the desorbed species approaches the material thickness $t_{m}$. They have a dependence with temperature since both depend on the coefficient of diffusion $D_{bulk}$, which is given by
$D_{\text{bulk}}=D_{\text{bulk,0}}\cdot\exp{\left(-\frac{E^{*}}{k\,T}\right)},$
where $E^{*}$ represents the activation energy, $k$ is the Boltzmann constant, and $T$ is the temperature. The time constant $\tau$ is computed as $\tau=\frac{t_{m}^{2}}{\pi^{2}D_{\text{bulk}}}$. A typical value for technical plastics, obtained in an outgassing setup, is $D_{\text{bulk}}=(3.0\pm 1.2)\times 10^{-9}$ cm2 s-1 [38].666 Diffusion coefficients, $D_{\text{bulk}}$, larger by up to $1000$ times and $10$ times can be found through direct measurements (e.g., [39]) for Teflon and PMMA, respectively. Similarly, the low-concentration diffusion coefficient of N2 in PEEK [40] is about $10$ times larger than the one used in text. Such high diffusion coefficients, if applicable to outgassing, would result in Teflon being depleted within $\tau\sim\text{minutes}$ (eq. 2.7-bottom), which contradicts common experience. The diffusion coefficient assumed in this work, however, does not cause depletion of the outgassing species over the time scales discussed (several days), when considering practical thicknesses above 2 mm. The potential presence of depletion effects at time scales shorter than the ones considered here represents an optimistic scenario, that needs to be addressed experimentally, and whose discussion we opted for leaving out of this work.
@@
Outgassing rates are typically measured experimentally in vacuum and can be parameterized through the value obtained 10 hours after pumping ($Q_{\text{10h}}$) and the exponent $a$ [41] (the latter assumed to be 0.5 hereafter). For instance, in the case of Teflon at room temperature, $Q_{\text{10h}}$ typically ranges from $3.5\times 10^{-8}$ to $1\times 10^{-7}$ $\text{mbar}\cdot\text{l}\cdot\text{cm}^{-2}\cdot\text{s}^{-1}$ (for a recent review, see [42]). Mass spectroscopy indicates that the O2 content in outgassing is generally less than 10% of the total, while N2 and H2O concentrations vary considerably, ranging from nearly 70% H2O and 30% N2/CO to almost 100% N2 and 0% H2O [21, 41, 43]. For both H2O and N2, we consider the most unfavourable case (i.e., that they constitute 100% of the outgassing), together with the middle (‘typical’) and lowest (‘optimistic’) $Q_{\text{10h}}$ values reported for total outgassing rates in [42]. For O2, we consider a range between the value directly measured in [21] (‘optimistic’) and 10% of the middle value of $Q_{\text{10h}}$ reported in [42] (‘typical’). Since outgassing rates depend on material quality and surface finish, our approach has been to select a range that includes both typical outgassing values and the lowest ones found in the literature. Measurements in [21] are particularly relevant as they do not involve bake-out, thereby matching the expected experimental conditions under which large-volume TPCs are commissioned.
@@
Outgassing rates for PMMA and PEEK are found in the range $Q_{\text{10h}}=10^{-7}$$10^{-6}$ $\frac{\textnormal{mbar}\cdot{\textnormal{l}}}{\textnormal{cm}^{2}\cdot{\textnormal{s}}}$. For PMMA, these rates tend to be dominated by H2O, with no trace of O2 or N2 being detected (e.g., [21]). For PEEK, the situation is similar; however, the amount of O2 reported in the literature is at the %-level (e.g., [44]). As PMMA is typically used as a window for light detection [45] and PEEK is sometimes present in readout frames [46], one can anticipate that the use of these materials would be restricted to the end-caps, constituting at most 20% of the total outgassing surface in our default geometry. Thus, the contribution of PMMA and PEEK to outgassing can generally be made subdominant in the presence of a Teflon reflector, if these materials are selected from the lower end of the range of measured values. A compilation of the outgassing rates considered in this work can be found in table 2 (column 2).
@@
In the following, we will adopt the outgassing law in eq. 2.7 for studying the impact of H2O and N2 on TPC performance. For O2 (a priori the most critical case), we will perform the calculations assuming a more extreme scenario in which the outgassing remains constant at $Q=Q_{\text{10h}}$. In all cases, Teflon will be assumed to dominate the system outgassing.
@@
$D_{\text{x, He/Ne/Ar/Xe}}=\frac{0.00143\cdot T^{1.75}}{P\sqrt{\frac{2}{\frac{1}{M_{\text{x}}}+\frac{1}{M_{\text{He/Ne/Ar/Xe}}}}}\cdot[\sqrt[3]{V_{\text{x}}}+\sqrt[3]{V_{\text{He/Ne/Ar/Xe}}}]^{2}},$ (2.8)
@@
$\frac{P}{D_{\text{x},B_{1}+B_{2}}}=\frac{P_{B_{1}}}{D_{\text{x},B_{1}}}+\frac{P_{B_{2}}}{D_{\text{x},B_{2}}}.$ (2.9)
@@
It should be noted that additional mechanisms at the material-gas interface could, in principle, introduce a pressure dependence and an outgassing law different from the one assumed in vacuum. This would need to be elucidated through direct measurements. As with other assumptions employed earlier in the text, the aim of our study is to provide upper limits to the impact of impurities in an actual system and verify that they can be kept below potentially harmful levels through reasonable design choices. In this regard, there is no obvious mechanism by which outgassing could be enhanced as pressure increases, compared to the model proposed here.
@@ 2.3 Gas distribution scheme
Our baseline proposal for the gas distribution is based on axial rods (perforated plastic tubes) attached to the field-cage, with inlets positioned at the bottom-half and outlets at the top-half, as shown in Fig. 4 (top-left). The rationale for this design is to enable a straightforward implementation, as the field-cage is generally the least instrumented part of a TPC. This concept does not interfere with the end-cap design and, with appropriate choices, can ensure uniform gas distribution along the drift/axial direction. Sixteen equispaced axial rods run along the rectangular section of the TPC (eight for injection and eight for extraction) as shown in Fig. 4 (top-right), providing an inter-rod distance of nearly 1 m between centres. As demonstrated later, this arrangement is sufficient to efficiently sweep outgassing from the field-cage walls. The distributor rods are periodically perforated with a pattern that is discussed in the next section.
@@
Assuming that gas is continuously recirculated and that clean gas is injected (as discussed previously), and in the absence of gas leaks, the ultimate purity levels that can be achieved are limited by the capacity of the recirculation compressors. For reference, the ALICE TPC circulates gas at 20 m3/h (1 bar) while the NEXT-100 experiment is designed to circulate at approximately 5 m3/h (10 bar), [49]. Based on this, we will consider two different scenarios for a 10 bar system, referring to them as ‘typical’ (30% above NEXT flow: 6.5 m3/h) and ‘optimistic’ (20 m3/h), as summarised later in table 3.777We will not refer to normal litres in text. All units of volume must be understood at the given pressure.
@@ 3.1 Optimization of the distributor rods
A qualitative discussion of the main features of gas distributors has been left to appendix (section A). Two reference cases are considered: a capped distributor and an open one. In general grounds, an open distributor is less efficient, given that only a fraction of the gas exits through the holes and into the chamber –as little as a few % for high gas flow. It also leads to a higher pressure drop along the distributor due to friction, significantly greater directionality of the outgoing gas and a non-uniform discharge pattern. Low efficiency in combination with a directional and non-uniform discharge pattern will adversely affect the distribution of impurities throughout the chamber and the time required to achieve uniform mixing with the main gas. Therefore, we opted for a capped distributor, ensuring uniform and normal injection of all the circulated gas.
A large number of holes per gas distributor, on the order of several hundreds, has been proposed in previous large-volume TPCs (e.g., [10]). Optimising such a geometry, with each hole potentially down to $\mathcal{O}($mm) in size and each distributor rod extending over a length of 555~{}5m, is impractical when using numerical simulations. As an alternative, we opted for a parametric approach based on the balance equations for each hole section, which allows the solution to be obtained iteratively. We followed the classical formulation of [52], whereby: i) a discharge coefficient $C_{d}$ (between 0 and 1) regulates the loss of flow into the distributor holes,  — mass balance; ii) a regain coefficient $C_{r}$ (between 0 and 2) regulates the pressure increase along the distributor after the ensuing velocity drop,  — momentum balance (Fig. 15-right, section A of appendix); iii) an angular coefficient $\gamma$ regulates the angle of discharge relative to the normal:
$D^{2}(v_{\text{u}}-v_{\text{d}})$ $=$ $d^{2}C_{d}\sqrt{\frac{\Delta{P_{\text{hole}}}}{\rho/2}},$ (3.1)
$P_{\text{d}}-P_{\text{u}}$ $=$ $C_{r}\frac{1}{2}\rho(v_{\text{u}}^{2}-v_{\text{d}}^{2}),$ (3.2)
$\theta$ $=$ $\gamma\cdot\arctan\left(\sqrt{\frac{\frac{1}{2}\rho v_{\text{u}}^{2}}{\Delta{P}_{\text{hole}}}}\right).$ (3.3)
@@
$\Delta{P_{\text{hole}}}=\frac{P_{\text{u}}+P_{\text{d}}}{2}-P_{{}_{\text{TPC}}}.$ (3.4)
Losses due to friction will occur when the gas transits between adjacent holes; however, in the conditions discussed here, the effect is negligible. For each distributor topology, three main quantities have been evaluated through the iterative solution of the balance equations: the pressure difference between the distributor and the TPC at the position of each hole ($\Delta{P_{\text{hole}}}$), the velocity of the gas as it flows through it ($v_{\text{hole}}$), and its maximum discharge angle relative to the normal ($\theta_{\text{max}}$).
In the conditions studied here, as $d\ll{D}$ and the discharge velocity is both small and very similar for all holes, it is possible to approximate the above magnitudes by (see appendix, section B)):
$\Delta{P}_{\text{hole}}$ $\simeq$ $\frac{1}{2}\frac{1}{C_{d}^{2}}\rho v_{\text{hole}}^{2},$ (3.5)
$v_{\text{hole}}$ $\simeq$ $\frac{\Phi}{\textnormal{t.h.a.}},$ (3.6)
$\theta_{\text{max}}$ $\simeq$ $\gamma C_{d}\frac{d^{2}}{D^{2}}\frac{L_{\text{dist}}}{\Delta{L}},$ (3.7)
@@
In order to come to some plausible distributor geometries, the following design constraints were introduced: i) a diameter hole not much smaller than 1 mm, ii) a diameter of the distributor not much larger than 5 cm, iii) an inter-hole distance not larger than 5 cm, iv) a pressure drop below 10 mbar, v) a discharge velocity well below the speed of sound, and vi) a maximum discharge angle below or around $5^{\circ}$. Based on these requirements, the main challenge is to obtain a small discharge angle without neither letting the inter-hole distance and tube diameter grow too big, nor the hole diameter become too small. Four exemplary geometries satisfying the above conditions ($A$-$D$) are compiled in table 3, although many more combinations are obviously possible.
It might be argued that small distributor diameters like those given in $C$ and $D$ should be preferred, so as not to reduce the fiducial volume due to sheer penetration in the active volume or charging-up. However, for hole velocities above 1-2 m/s (the velocity present in geometries $C$ and $D$) we observed a tendency of the flow to form large-scale vortices. Those vortices substantially delayed the time for impurities to achieve uniform mixing. Aiming at a comparative study, we opted for a design that guaranteed a range of velocities in the default distributor geometry outside the region where the effect appears, for the gas flows studied. Furthermore, even for geometries $A$ and $B$, small distributor diameters can also be used through minor adjustments of the geometry or increasing the thickness of the distributor walls (studies that fall outside the scope of this work) [^8] For instance, a reduction from a diameter of 7 cm to 3.5 cm in geometry $B$ can be achieved by increasing the number of injection rods to 32 and increasing the inter-hole distance to 60 mm, yet keeping the rest of the fluid properties as in table 3.. Therefore, we have focused in the remainder of the text on $A$ (single hole) and $B$ (triple hole), taking the latter as the default distributor geometry. We wish to recall the general wisdom that field distortions in a field cage decrease quickly when the distance to the feature (distributor, in this case) extends by ×$\times 2$ its size, and uniform charging-up (a plausible assumption under uniform irradiation or under the use of electrostatic-dissipation transparent coatings [53]) will theoretically reduce the extent of field distortions to zero. As a consequence, it is expected that, even in a worst-case scenario, a generous fiducialisation of up to 10-20 cm (radially inwards) will suppress most (if not all) of the effect stemming from charging-up of the Teflon and distributors, in geometries $B$ and $D$. We plan to discuss this as part of a separate paper, but details are too context-dependent and out of scope.
A CAD image corresponding to geometry $B$ is shown in Fig. 5-a, whereas Fig. 5-b provides an equivalent geometry in terms of fluid parameters at the inlets, however with smaller-diameter distributors.
@@
An example of the results obtained is presented in Fig. 6 (for details on the numerical procedure, the reader is referred to the appendix –section A). The analytical estimates detailed previously are included as blue dashed lines, showing excellent agreement with the numerical solution. It should be noted that, with independence from the distributor geometry, $C_{d}$ approaches a constant value around $C_{d}=0.61$-$0.64$ when $\Delta{P}_{\text{hole}}\gg\frac{1}{2}\rho v_{\text{dist}}^{2}$ [54]. A detailed comparison with experimental data is presented in [52], providing an estimate of $C_{d}=0.63$ (which is adopted hereafter). Such pressure-dominated conditions are typical of capped distributors and were verified for all our study cases. Similarly, the $\gamma$ pre-factor is largely independent of the gas flow characteristics and, according to [52], weakly dependent on the geometry, and so the value is set to $\gamma=0.71$. The least constrained coefficient, $C_{r}$, as well as the friction losses, are immaterial in the conditions discussed, as is made clear through the analytical expressions in 3.5-3.7. The aforementioned values for the parameters $\gamma$ and $C_{d}$ are consistent with the CFD analysis of the control geometry displayed in Fig. 15 (appendix).
Figure 6: Main magnitudes characterizing the response of the distributor rod for each triple-hole along the distributor, shown with red lines (the first triple-hole upstream is number 1). We consider an argon gas flow of 6.5 m3/h at 10 bar, 8 distributors of 3 cm in diameter and 5 m in length. The hole size is 0.75 mm and holes are separated by 5 cm (geometry $D$ in table 3). Top-left: gas velocity in the distributor at the hole position. Top-right: discharge velocity. Bottom-left: angle relative to normal. Bottom-right: pressure difference relative to the TPC. (The analytical expressions given in text are shown as dashed lines).